OpenStudy (anonymous):

x2+12x-4=0 complete the square

OpenStudy (inkyvoyd):

Heey. Wanna read a very long tutorial on how to complete the square?

OpenStudy (inkyvoyd):

3 days ago, I spent 4 hours writing it :)

OpenStudy (anonymous):

x2+12x +36 -40=0

OpenStudy (inkyvoyd):

Or you can give myko a medal -.-

OpenStudy (inkyvoyd):

btw x2+12x +36 -40=0 => x^2+2*6x+6^2-40=0